Joey’s Bike Shop sells new and used bicycle parts. Although a majority of its sales are cash sales, it makes a significant amount of credit sales. During 2018, its first year of operations, Joey’s Bike Shop experienced the following:

Sales on account $ 260,800
Cash sales 634,400
Collections of accounts receivable 247,760
Uncollectible accounts charged off during the year 1,096

Required

Assume that Joey's Bike Shop uses the allowance method of accounting for uncollectible accounts and estimates that 1 percent of its sales on account will not be collected. Answer the following questions:

  1. What is the Accounts Receivable balance at December 31, 2018?
  2. What is the ending balance of Allowance for Doubtful Accounts at December 31, 2018, after all entries and adjusting entries are posted?
  3. What is the amount of uncollectible accounts expense for 2018?
  4. What is the net realizable value of accounts receivable at December 31, 2018?
a. Accounts receivable
b. Allowance for doubtful accounts
c. Uncollectible accounts expense
d. Net realizable value
Answer #1

Accounts receivable = Sales on account - Collections of accounts receivable - Uncollectible accounts charged off during the year

= 260,800 - 247,760 - 1,096

= $11,944

Ending balance of Allowance for doubtful accounts = Uncollectible accounts expense -  Uncollectible accounts charged off during the year

= 2,608 - 1,096

= $1,512

Uncollectible accounts expense = 1 percent of sales on account

= 260,800 x 1%

= $2,608

Net realizable value of Accounts receivable = Accounts receivable - Ending balance of Allowance for doubtful accounts

= 11,944 - 1,512

= $10,432

a. Accounts receivable = $11,944

b. Allowance for doubtful accounts = $1,512

c. Uncollectible accounts expense = $2,608

d. Net realizable value = $10,432